# Learning from History: Task-agnostic Model Contrastive Learning for Image Restoration

### 18 Dec 2023 v2

@@ Introduction
We first investigate the image super-resolution task, utilizing both CNN-based EDSR and Transformer-based SwinIR as baselines to evaluate our proposed self-pprior guided negative regularization. Results presented in Tab. 1 across various architectures and model capacities. Our proposed model contrastive paradigm can achieve considerable improvements when retraining lightweight models such as EDSR and SwinIR, achieving average gains of 0.07 dB and 0.09 dB on scale $\times$4, respectively. Moreover, the retrained EDSR model demonstrats a noteworthy average improvement of 1.3 dB for the scale $\times$2 task. Notably, there are substantial improvements by our retrained on the Urban100 and Manga109 datasets. This implies that our model contrastive paradigm facilitates a more accurate and compact convergence in the restoration process. Moreover, comparisons with current state-of-the-art methods are presented in Fig. 2 (a). Remarkably, the retrained EDSR and SwinIR can eclipse the performance of several contemporary methods. These underscore the promise of our model contrastive paradigm.
In addition to the lightweight models, we also retrain the large SwinIR model. Our proposed approach exhibits its general applicability to larger models, providing an average improvement of 0.05 dB. It is worth noting that the improvements seen in larger models were relatively smaller when compared to the lightweight models. This observation is consistent with the understanding that lightweight models, owing to their smaller capacity, may struggle to find an optimal solution. Therefore, our proposed model contrastive paradigm provides significant assistance in these cases. Conversely, larger models have the capacity to fit or even overfit the dataset, hence the model contrastive paradigm improves them within a lesser margin. Visual results are showcased in Fig. 4. It is also evident that our model contrastive paradigm can improve existing methods, particularly in bringing more precise textures.
@@
We take the state-of-the-art IDT (Xiao et al. 2022) as the baseline and retrain it by the proposed model contrastive paradigm on five datasets ,and Tab. 2 provides a comprehensive comparison. The retrained IDT, achieves an average PSNR improvement of 0.26 dB across all test datasets. For instance, it showcase a gain of 0.7 dB in comparison to its original on SPA dataset. Furthermore, visual results are illustrated in Figure 5, accompanied by a divergence map. It intuitively highlights the enhancements in the degraded regions. These insights underscore the capability of our model contrastive paradigm to refine the target degradation.
@@
Table 3: Comparison results of image dehazing. Reults of our retrained models are in bold.
Method MIMO-UNet HINet MAXIM Restormer UFormer NAFNet NAFNet+SPN (Ours)
PSNR 32.68 32.71 32.86 32.92 32.97 32.87 32.93 (+0.06)
SSIM 0.959 0.959 0.961 0.961 0.967 0.9606 0.9619 (+0.0013)
@@
Following the recent work (Song et al. 2023), we utilize the SOTS-indoor and SOTS-mix datasets for testing. In detail, we retrain FFANet (Qin et al. 2020b) and DehazeFormer (Song et al. 2023), with the outcomes detailed in Fig. 3. From the table, a remarkable enhancement is achieved over FFANet, exhibiting gains of 3.41 dB and 0.69 dB on the indoor and mixed test datasets, respectively. Concurrently, the Transformer-based model IDT also record advancements, evident across various model scales. Specifically, the retrained DehazeFormer-B achieves the most pronounced improvement of 0.57 dB on the indoor test dataset. Intriguingly, the retrained FFANet even outperformes the DehazeFormer. It is reasonable that FFANet is a CNN-based model, which has a large model capacity with heavy parameter counts, and our model contrastive paradigm appears to steer FFANet towards more optimized results. Some visual results are presented in Fig. 6. One can find that retrained FFANet is clearer with less artifacts.
@@
FFANet (Qin et al. 2020b) 36.39 0.9886
+CR (Wu et al. 2021) Image Dehazing 36.74 0.9906
+CCR (Zheng et al. 2023) (SOTS-indoor ) 39.24 0.9937
@@
EDSR (Lim et al. 2017) SISR 26.04 0.7849
+PCL (Wu, Jiang, and Liu 2023) (Urban100) 26.07 0.7863
@@
Table 5: A comparative analysis of the existing contrastive paradigms versus our proposed model contrastive approach. Existing methods are task-oriented and proposed for image dehazing and image super-resolution separately. Our model contrastive paradigm is task-agonist and general to improve diverse image restoration tasks.
Comparison to Existing Sample Contrastive Paradigm
In contrast to existing methods that employ task-oriented negatives, our model contrastive paradigm provides straightforward enhancement to image restoration tasks. We furnish a comprehensive comparison between exiting contrastive approaches and our innovative model contrastive paradigm. For image dehazing task, we examine previous methods based on FFANet, including contrastive regularization (CR) (Wu et al. 2021), curricular contrastive regularization (CCR) (Zheng et al. 2023), and our self-prior guided negative loss (SPN). The comparative results can be found in Tab. 5. Instead of merely employing low-quality input images as negative samples, CCR achieves significant advancements by utilizing consensus negatives from pre-trained models and the curriculum learning approach. Intriguingly, our SPN outperforms other methods, achieving the best improvement without any additional prior knowledge. In addition, for image super-resolution (Wu, Jiang, and Liu 2023), our model contrastive paradigm again registers the highest gains, as presented in Tab. 5. This consistent performance underscores the fact that our proposed model contrastive paradigm is task-agnostic and highly effective for diverse image restoration tasks.
@@
To provide a deeper understanding for the role of our negative model, our ablation study involves the exploration of different configurations of the negative model, including an randomly initialized and fixed SwinIR model, a pre-trained and fixed SwinIR model, and our default one. The results of these different configurations, which shed light on the efficacy of our proposed approach, are presented in Tab. 6. ’-’ presents the baseline model without contrastive negative model. Firstly, negative model with random and fixed parameters can not afford effective negatives and impedes the target SR model. When we set the negative model as the pre-trained and fixed one, it enhances the target model but achieves worse performance than ours. We think the adaptive and curricular negatives generated by model contrastive paradigm is essential to provide a smooth negative regularization, which is consistent with some existing methods.
@@
We conducted an ablation study, where FFANet was retrained with various negative models, each differentiated by its updating iteration, $s$. The results are tabulated in Tab. 7. It becomes evident that the choice of step $s$ distinctly influences both the negative model and performance. A smaller step provides more challenging negatives compared to the larger step, leading to superior performance. In our experiments, we employ multiple negative models with the four steps together, achieving the best results. The incorporation of multiple negative models ensures a consistent supply of robust negatives.
@@
We study the influence of the coefficient $\lambda$ in Eq. (1), and results are presented in Tab. 8, where 0 value means the original $L_{1}$ loss without our negative regularization. When $\lambda$ set as 1e-2 or 1e-3, there is collapse during training, especially in the early stage. This is reasonable that large negative loss can influence the model updating towards to the optimal direction. When $\lambda$ is 1e-4, our model contrastive paradigm achieves the best performance. Considering that the proposed model contrastive paradigm is task-agonist and adaptive to various image restoration tasks, in this paper, we take $\lambda=1e-4$ as the default value across different tasks.
$\lambda$ Value 0 1e-2 1e-3 5e-4 1e-4 1e-5
@@
In reference to Eq. (1), we examine the influence of varying updating weights. The outcomes are detailed in Tab. 9. In general, a smaller value of $w$ retains less latency information, bringing it in closer alignment with the target model. This results in the generation of the most challenging negatives. The symbol ’-’ indicates model collapse during training. As $w$ increases, it becomes evident that the optimal performance is attained at $w=0.1$. Based on our empirical observations, we consistently utilize $w=0.1$ as the default setting in all subsequent experiments.
@@
In this paper, we proposed a task-agnostic model contrastive paradigm for low-level tasks, improving existing models. While our experiments cover a range of tasks, areas like JPEG artifact removal, image denoising, and certain real-world scenarios remain unexplored. Our ablation studies focused on image super-resolution, as shown in Tables 6 and 7, but we did not extensively evaluate our hyper-parameters across all tasks. These unexamined aspects offer valuable opportunities for future research.
@@ Conclusion
In this study, we propose innovative model contrastive paradigm for various low-level tasks. Compared to the task-oriented negatives in existing methods, the proposed model contrastive paradigm, constructing negatives from the target model itself, a is task-agonist, general to various image restoration tasks by a Self-Prior guided Negative loss (SPN). Our proposed SPN is straightforward to implement and does not rely on any task or data prior. We have retrained several image restoration models, and they achieve significant improvements across various tasks and architectures. In the future, we believe it would be meaningful to evaluate our proposed paradigm in more dense prediction tasks, potentially offering fresh insights and advances for the community.
Acknowledgments The research was supported by the National Natural Science Foundation of China (U23B2009, 92270116).

### 19 Jan 2024 v4

@@ Methods
We first investigate the image super-resolution task, utilizing both CNN-based EDSR and Transformer-based SwinIR as baselines to evaluate our proposed self-prior guided negative regularization. Results presented in Tab. 1 across various architectures and model capacities. Our proposed model contrastive paradigm can achieve considerable improvements when retraining lightweight models such as EDSR and SwinIR, achieving average gains of 0.07 dB and 0.09 dB on scale $\times$4, respectively. Moreover, the retrained EDSR model demonstrates a noteworthy average improvement of 1.3 dB for the scale $\times$2 task. In particular, there are substantial improvements from our re-training on the Urban100 and Manga109 datasets. This implies that our model contrastive paradigm facilitates a more accurate and compact convergence in the restoration process. Moreover, comparisons with current state-of-the-art methods are presented in Fig. 2 (a). Remarkably, the retrained EDSR and SwinIR can eclipse the performance of several contemporary methods. These underscore the promise of our model contrastive paradigm.
In addition to the lightweight models, we also retrain the large SwinIR model. Our proposed approach exhibits its general applicability to larger models, providing an average improvement of 0.05 dB. It is worth noting that the improvements seen in the larger models were relatively smaller compared to the lightweight models. This observation is consistent with the understanding that lightweight models, owing to their smaller capacity, may struggle to find an optimal solution. Therefore, our proposed model contrastive paradigm provides significant assistance in these cases. Conversely, larger models have the capacity to fit or even overfit the dataset, hence the model contrastive paradigm improves them within a lesser margin. Visual results are showcased in Fig. 3. It is also evident that our model contrastive paradigm can improve existing methods, particularly in bringing more precise textures.
@@
We take the advanced IDT (Xiao et al. 2022) as a baseline and retrain it by the proposed model contrastive paradigm on five datasets, and Tab. 2 provides a comprehensive comparison. The retrained IDT achieves an average PSNR improvement of 0.26 dB across all test datasets. For instance, it showcases a gain of 0.7 dB in comparison to its original in the SPA dataset. Furthermore, visual results are illustrated in Figure 4, accompanied by a divergence map. It intuitively highlights enhancements in the degraded regions. These insights underscore the ability of our model contrastive paradigm to refine target degradation.
Table 3: Comparison results of image dehazing. The results of our retrained models are in bold
@@
Table 3: Comparison results of image dehazing. The results of our retrained models are in bold
@@
Following the recent work (Song et al. 2023), we utilize the SOTS-indoor and SOTS-mix datasets for testing. In detail, we retrain FFANet (Qin et al. 2020b) and DehazeFormer (Song et al. 2023), with the results detailed in Fig. 3. From the table, a remarkable enhancement is achieved over FFANet, exhibiting gains of 3.41 dB and 0.69 dB on the indoor and mixed test datasets, respectively. Concurrently, the Transformer-based model DehazeFormer also records advancements, evident across various model scales. Specifically, the retrained DehazeFormer-B achieves the most pronounced improvement of 0.57 dB on the indoor test dataset. Intriguingly, the retrained FFANet even outperforms the DehazeFormer. It is reasonable that FFANet is a CNN-based model, which has a large model capacity with heavy parameter counts, and our model contrastive paradigm appears to steer FFANet towards more optimized results. Some visual results are presented in Fig. 5. One can find that the retrained FFANet is clearer with fewer artifacts.
@@
Comparison to Existing Sample Contrastive Paradigm
In contrast to existing methods that employ task-oriented negatives, our model contrastive paradigm provides straightforward enhancement to image restoration tasks. We furnish a comprehensive comparison between exiting contrastive approaches and our innovative model contrastive paradigm. For the image dehazing task, we examine previous methods based on FFANet, including contrastive regularization (CR) (Wu et al. 2021), curricular contrastive regularization (CCR) (Zheng et al. 2023), and our self-prior guided negative loss (SPN). The comparative results can be found in Tab. 5. Instead of merely employing low-quality input images as negative samples, CCR achieves significant advancements by utilizing consensus negatives from pre-trained models and the curriculum learning approach. Intriguingly, our SPN outperforms other methods, achieving the best improvement without additional prior knowledge. In addition, for image super-resolution (Wu, Jiang, and Liu 2023), our model contrastive paradigm again registers the highest gains, as presented in Tab. 5. This consistent performance underscores the fact that our proposed model contrastive paradigm is task-agnostic and highly effective for diverse image restoration tasks.
Methods MIMO-UNet HINet MAXIM Restormer UFormer NAFNet NAFNet+SPN (Ours)
PSNR 32.68 32.71 32.86 32.92 32.97 32.87 32.93
SSIM 0.959 0.959 0.961 0.961 0.967 0.9606 0.9619
@@
FFANet 36.39 0.9886
+CR Image Dehazing 36.74 0.9906
+CCR (SOTS-indoor ) 39.24 0.9937
@@
EDSR SISR 26.04 0.7849
+PCL (Urban100) 26.07 0.7863
@@
Table 5: A comparative analysis of the existing contrastive paradigms versus our proposed model contrastive approach. Typically, existing methods are task-oriented and are proposed for image dehazing (CR (Wu et al. 2021) and CCR (Zheng et al. 2023)) and image super-resolution (PCL (Wu, Jiang, and Liu 2023)), separately. Our model contrastive paradigm is task-agonist and outperform existing methods.
@@
This ablation study aims to elucidate the role of our negative model by investigating various configurations, including a randomly initialized and fixed SwinIR model, a pre-trained and fixed SwinIR model, and our standard model. The results, detailed in Tab. 6, highlight the effectiveness of our approach. The baseline model is denoted by ’-,’ indicating the origin SwinIR. We found that a negative model with random and fixed parameters hinders the target SR model, failing to provide effective negatives. Conversely, using a pre-trained and fixed negative model shows improvement, but not as much as our model contrastive. This suggests the importance of adaptive and curriculum-based negative generation in our model contrastive paradigm, aligned with the findings of existing methods (Zheng et al. 2023).
@@
We conducted an ablation study, in which FFANet was retrained with various negative models, each differentiated by its update iteration, $s$. The results are tabulated in Tab. 7. It becomes evident that the choice of step $s$ distinctly influences both the negative model and performance. A smaller step provides more challenging negatives compared to the larger step, leading to superior performance. In our experiments, we employ multiple negative models with the four steps together, achieving the best results. The incorporation of multiple negative models ensures a consistent supply of robust negatives.
@@
We study the influence of the coefficient $\lambda$ in Eq. (3), and results are presented in Tab. 8, where the value 0 means the original loss $L_{1}$ without our negative regularization. When $\lambda$ is set to 1e-2 or 1e-3, there is collapse during training, especially in the early stage. This is reasonable because large negative loss can influence the model updating towards the optimal direction. When $\lambda$ is 1e-4, our model contrastive paradigm achieves the best performance. Considering that the proposed model contrastive paradigm is general to various image restoration tasks, in this paper we take $\lambda=1e-4$ as the default value across different tasks.
$\lambda$ 0 1e-2 1e-3 5e-4 1e-4 1e-5
@@
In reference to Eq. (3), we examine the influence of varying updating weights. The outcomes are detailed in Tab. 9. In general, a smaller value of $w$ retains less latency information, bringing it in closer alignment with the target model. This results in the generation of the most challenging negatives. The symbol ’-’ indicates model collapse during training. As $w$ increases, it becomes evident that the optimal performance is attained at $w=0.1$. On the basis of our empirical observations, we consistently use $w=0.1$ as the default setting in all subsequent experiments.
@@
In this paper, we proposed a novel model contrastive paradigm for low-level image restoration tasks, improving existing models. Although our experiments cover a range of tasks, areas such as JPEG artifact removal, image denoising, and certain real-world scenarios remain unexplored. Our ablation studies focused on image super-resolution, as shown in Tab. 6 and  7 but we did not extensively evaluate our hyperparameters across all tasks. These unexamined aspects offer valuable opportunities for future research.
@@ Conclusion
In this study, we propose an innovative model contrastive paradigm for various low-level tasks. Compared to the task-oriented negatives in existing methods, the proposed model contrastive paradigm, constructing negatives from the target model itself, is task-agonist and general to various image restoration tasks by a self-prior guided negative loss (SPN). Our proposed SPN is straightforward to implement. We have retrained several image restoration models, and they achieve significant improvements across various tasks and architectures. In the future, we believe it would be meaningful to evaluate our proposed paradigm in more dense prediction tasks, potentially offering fresh insights and advances for the community.
Acknowledgments
The research was supported by the National Natural Science Foundation of China (U23B2009, 92270116).